OpenStudy (anonymous):

if a species contains 23% A in its DNA what is the percentage of guanine it would contain? A. 23% B. 46% c.25% d.44% e.27%

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

ah really good explanation and well understood thank you very much

OpenStudy (anonymous):

Chargaff-rule states: A + G = T + C = 50%. 50 - 23 = 27. :-)